Two objects P and ! carry charges -3.00µC and -0.44µC. What is the number of electrons transferred from P to Q, so that they acquire equal charges?

Grade:12

2 Answers

Ayush Jain
6 years ago
Ans. 8x1012
Calculate the charge on each object by 1/2(q1+q2)
then calculate difference in charge and divide it by charge on electron.
Naman Jain
6 years ago
Charge Needed on each object to be equal = ½ [(-3)+(-0.44)]  microCoulomb
=> ½ × (-3.44) = -1.72 microCoulomb
Q is already have the charge -0.44 mC itself,
so it needs [(-1.72)-(-0.44)] = -1.28mC charge from P to be equal
 
Since,
Q = ne
n = Q/e
=> -1.28×10^-6
       1.6×10^-19
= 0.8×10^13
=> n = 8 × 10^12